LUCI prep/flexure testing:
UT 2011 Mar 07

Observer: DThompson @LBT, OKuhn, ARakich in Tucson
Telescope Operator: GBechetti
Instrument Support: DOfficer
Telescope Support: MMidkiff
SW support: SHooper?

Summary

Preparations

Details

LBCFPIA work: We (A Rakich and O Kuhn) changed three values in the FPIA code in /home/lbcobs/LBCFPIA/lbcfpia/src tonight:
  • in lbcfpia.cfg:
    • M12 element of Blue Matrix, from -2.4 to -1.75. This value was determined from tests in which we ran dofpia with a stringent convergence criterion, and when the usual convergence criteria are met, we drove in +/- 1000nm Z11. We measured the change in pupil diameter in the resulting pupil to be -/+ 5 pixels. The Z4 needed to correct this is -/+1750nm. i.e., after driving in +1000nm Z11, FPIA wants to send -1000nm, and it should also send +1750nm Z4. Now the matrix is:
      BlueZIMatrix      = [[1.0,-1.75,-8.0],[0.0,1.0,3.0],[0.0,0.0,1.0]] 
    • BlueZ22Zero from 2.3 to 1.8 (recent changes to this parameter have been, from 1.6 to 2.3, in Dec 2010, and now from 2.3 to 1.8. Note that a filter focus offet must be re-measured with this new value
        BlueZ22Zero      = 1.8
  • in sendsph3.pro
    • The term to relate z4 and z11 was changed from -1.4 to -1.75 to match the value of the blue M12 matrix element.
   res[3] = -1.75 * res[10]
  • I also made changes to the program calcabcoeffs.pro so that in dbg mode, Z22 aberrations would also be computed. These three new lines are:
LogAllWays,"Computed ab's: "+'Z4 = '+strtrim(ab[3],2)+ '  Z5 = '+ strtrim(ab[4],2)+$
           '  Z6 = '+strtrim(ab[5],2)+ '  Z7 = '+ strtrim(ab[6],2)+$
           '  Z8 = '+strtrim(ab[7],2)+ '  Z11 = '+strtrim(ab[10],2)+$
           '  Z22 = '+strtrim(ab[21],2)
LogAllWays,"xtalk corr'd : "+'Z4 = '+strtrim(ab[3],2)+ '  Z5 = '+ strtrim(ab[4],2)+$
           '  Z6 = '+strtrim(ab[5],2)+ '  Z7 = '+ strtrim(ab[6],2)+$
           '  Z8 = '+strtrim(ab[7],2)+ '  Z11 = '+strtrim(ab[10],2)+$
           '  Z22 = '+strtrim(ab[21],2)
LogAllWays,"Gain corr'd  : "+'Z4 = '+strtrim(ab[3],2)+ '  Z5 = '+ strtrim(ab[4],2)+$
           '  Z6 = '+strtrim(ab[5],2)+ '  Z7 = '+ strtrim(ab[6],2)+$
           '  Z8 = '+strtrim(ab[7],2)+ '  Z11 = '+strtrim(ab[10],2)+$
           '  Z22 = '+strtrim(ab[21],2)

  • To be done (first item can be done if we are able to open tonight; rest of the tests to be done tomorrow or the next night):
    • Check the Z4/Z11 relation now that the abovementioned changes have been made.
      • Run dofpia with stringent limit criteria: dofpia, BlueLimit=10, RedLimit=10
      • When convergence to the usual limits is reached: (z4,z5,z6, z11 <~ 250nm, z7,z8 <~ 400nm and z22 <~50nm) and the converged pupil image is reading out, send Z11 = +/- 1000nm. Note the before/after image numbers and allow a few more iterations, until it appears reasonably well-converged.
    • Blue filter focus offsets need to be remeasured after this change to BlueZ22Zero.
    • Redetermine the matrix elements M13 (Z4/Z22; currently 8.0) and M23 (Z11/Z22, currently 3.0):
      • Follow above steps to check Z4/Z11 relation, but when the converged pupil image is reading out, send Z22 = +/- 200nm.
    • Astigmatism has appeared to be oscillating in recent logs. Change AstigScale by about 60% (from 220 to 130) and test this change.

Observing Plan (copy and paste this to the nightly logs)

LUCI1 functional checkout:
  • Check system time vs our GPS, at cmd line execute: "ntpdate -q 10.144.0.211" OK
  • Run /home/engineer/bin/showSupervisorStatus.sh OK
  • Check LUCI1 Web-IO Control, verify that ROE and MCE are powered ON
  • Start LUCIFER Control SW and check that all is running as needed
    • Restart TelescopeService if TCS has been restarted TEL and GEIRS restarted
  • Open UI and check that telescope data are being updated YES
  • Check darks (see: LUCIFER$/readnoise script)
    • O2DCR 2.0s (usually around 11.57 e-) Files 2,3: RN=11.59e-
    • MER10 10.0s (usually around 5.2 e-) Files 4,5: RN=5.45e-
  • Take sieve mask images in N3.75 and N1.8 to check focus gradients
  • Move calibration unit in already in
    • Check that lamps work YES
    • Move calibration unit out YES

  • Run LUCI1 flexure tests, see: ~/LBTO/Flexures/RotatorScript.txt
    • Files 0008-0036; plot attached below (mirror_N375_20110307.ps)
    • Files 0068-0097; plot attached below (mirror_N180_20110307.ps)
    • Files 0147-0176; plot attached below (mirror_N180_20110307.2.ps)
    • Files 0207-0236; plot attached below (mirror_N375_20110307.2.ps)
    • Info sent to Walter.

  • Check if FLUSH_DETECTOR now saves data because of autosaves
    • The FLUSH_DETECTOR command uses DIT=0 NDIT=1 NEXPO=150
    • With save image = manual on RMGUI no files saved
    • With save imnage = automatic on RMGUI no files saved

Note: On a fresh login to the observer desktop on lucifer, we now get a firefox and the GEIRS init window on desktop 1 and three xterms on desktop 2!

Note: UI was up at 4:30 when I arrived at the telescope. A logout and log back in to observer did not clear the GEIRS rsync, it was still trying to copy data from the previous night. Logout, restart GEIRS Server, and log back in cleared it.

Note: Still a problem with the rsync, called Stephen who noted many java errors in dmesg. At about 3:00UT, with Stephen on the phone we shutdown and restarted the LUCI SW.

-- DavidThompson - 07 Mar 2011
Topic revision: r5 - 09 Mar 2011, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback